Two 200 pounds lead ball are separated by the distance of 1 meter .both ball have sam positive charge q.what charge will produce in electrostatic force between the that is of the same order of magnitude as the weight of one ball.
1
Expert's answer
2020-06-04T09:56:52-0400

First, convert 200 pounds to kilograms: multiply 200 by 0.454: 90.7 kg.

Second, calculate the weight of the ball:


"P=mg."


On the other hand, this force must be equal to the force of electrostatic interaction:


"P=mg=\frac{1}{4\pi\epsilon_0}\frac{q^2}{r^2},\\
\space\\
q=2r\sqrt{mg\pi\epsilon_0}=2\cdot1\sqrt{90.7\cdot9.8\cdot3.14}=106\text{ C}."